OpenStudy (anonymous):

Solve: 4(x + 6) = 6(x − 2)

OpenStudy (geometry_hater):

Distribute the numbers outside the parentheses

OpenStudy (anonymous):

First distribute. 4x + 24 = 6x - 12 Now get x by itself.

OpenStudy (anonymous):

X=18

OpenStudy (anonymous):

4x+24=6x-12 4x-6x=-12-24 -2x=-36 Divide both sides by -2. x=18

OpenStudy (anonymous):

2x = 36 x = 18. Very good